OpenStudy (anonymous):

9bc/10c^5 x 2c/3b multiply and simplify as much as possible

OpenStudy (mathmath333):

\(\large \color{black}{\begin{align} & \dfrac{\dfrac{9bc}{10c^5 \times 2c}}{{3b}} \hspace{.33em}\\~\\ & =\dfrac{\dfrac{9b}{20c^5}}{{3b}} \hspace{.33em}\\~\\ & =\dfrac{3b}{20c^5 } \hspace{.33em}\\~\\ \end{align}}\)

OpenStudy (anonymous):

i got 3b^2c^7/5

OpenStudy (anonymous):

am i right?

OpenStudy (mathmath333):

first thing is u should write properly the question in brackets or latex is it \(\large \color{black}{\begin{align} & \dfrac{\dfrac{9bc}{10c^5 \times 2c}}{{3b}} \hspace{.33em}\\~\\ \end{align}}\) or \(\large \color{black}{\begin{align} & \dfrac{9bc}{10c^5} \times \dfrac{ 2c}{3b} \hspace{.33em}\\~\\ \end{align}}\)